OpenStudy (anonymous):

Determine the number of elements in the sample space for each of the following: a. Tossing a coin three times b. The order that the top 5 students will receive their diplomas. c. Tossing a pair of dice.

OpenStudy (anonymous):

2 possibilities for each coin, by the counting principle it is \[2\times 2\times 2=2^3\]

OpenStudy (anonymous):

c) six possibilities for each die, by the counting principle it is \(6\times 6=6^2\)

OpenStudy (anonymous):

b) 5 choices for the first place 4 choices for the second 3 for the third 2 for the fourth 1 for fifth counting principle say multiply these together get \[5\times 4\times 3\times 2\times 1=5!\]

OpenStudy (anonymous):

thnx